Kutztown University Department of Mathematics

Colloquium

3:30 p.m.

Tuesday, November 10, 2009

Boehm Hall 260

The Stable Marriage Problem

Dr. Ju Zhou

Assistant Professor of Mathematics Kutztown University of Pennsylvania

Abstract

In a matchmaking club, if there are n boys and n girls, is there a way to pair them to date? If yes, how many different ways can the matchmaker pair them to date? Furthermore, if each girl has a rank of preference of the boys and each boy has a rank of preference of the boys, is there a way to pair them to date and make sure the relationship stable?

In this talk, some basic definitions of graph theory will be introduced. Then the problems above will be addressed and the corresponding results will be introduced.

3:00 p.m. refreshments served

3:30 p.m. talk begins

Kutztown Univ. of Pennsylvania Dept. of Mathematics Dr. Paul S. Ache, III, Chair